KaliHood:

Paolo is putting a fence around his pool. The length of the pool is 20 feet, the width is (20 minus x) feet, and the area of the pool is 300 square feet. Recall that area is length multiplied by width. The equation below models the area of the pool. 20 (20 minus x) = 300 What is the value of x? Round to the units place.

Hoodmemes:

The lenth = 20 ft and width = 20-x The area=\[lenth \times width\]